what is the present value of the following future amount? $60,924, to be received 20 years from now, discounted back to the present at 5 percent, compounded annually.
Answer #1

PRESENT VALUE = FUTURE VALUE/(1+RATE)n

PRESENT VALUE = 60924/(1+0.05)20 = 60924/2.6532977 = 22961.61 (NO INTERMEDIATE ROUNDING IS DONE)

ANSWER : $22961.61 (THUMBS UP PLEASE)